OpenStudy (florida.girl):

Here is a little cool "magic trick" Think of a number any number in the world Now add 13 to your number Then add 5 Then subtract your original number that you got in the beginning You got 18 correct? Please comment if you got it!!

OpenStudy (amtran_bus):

145 145+13 = 158 158+5 = 163 163-145 = 18

OpenStudy (amtran_bus):

lol i mean all you are doing is adding 18 to the orig number and subtracting it back.

OpenStudy (florida.girl):

Yea lol a lot of people don't get that though lol😂😂

OpenStudy (dumbcow):

Here's a better one: Pick any number Add 10 Multiply by 2 Subtract 8 Divide by 2 Subtract original number You will always get 6.

ganeshie8 (ganeshie8):

Nice one! Did you invent this trick @Florida.Girl ?

undeadknight26 (undeadknight26):

I've seen it around but it's still very very clever!

OpenStudy (robtobey2):

\[n + 13 + 5 - n = 18 \]

OpenStudy (zzr0ck3r):

Here is a better one Think of a number between 1 and 10 multiply it by 2 now add 8 now divide by 2 Now subtract your original number Now think of the letter in the alphabet associated with your number. 1=a, 2=b, 3=c,.... Now think of a country that starts with that letter Now go to the next letter in the alphabet now think of a animal that starts with that letter here is the thing There are no elephants in Denmark

OpenStudy (florida.girl):

Nope @ganeshie8 a friend told it to me

zepdrix (zepdrix):

.